A) Your microwave oven is 64% efficient. How much electrical energy must it use to produce 900 J of radiant energy? Show all your work.


1
Expert's answer
2022-05-25T10:26:57-0400
"\eta=\frac{W}{E_{tot}}"

"0.64=\frac{\rm 900\: J}{E_{tot}}"

"E_{tot}=\rm\frac{900\: J}{0.64}=1406\: J"